Dissenting Views About Mathematics and Mathematicians

• “A mathematician is a blind man in a dark room looking for a black cat which isn’t there,” Charles Darwin

• “Mathematicians are like Frenchmen. Whatever you say to them, they translate into their own language, and forthwith it is something completely different,” Johann Wolfgang von Goethe (1749 - 1832, German author and philosopher)

• “As far as the laws of mathematics refer to reality, they are not certain; and as far as they are certain, they do not refer to reality,” Albert Einstein

• “Mathematicians, who are only mathematicians, have exact minds, provided all things are explained to them by means of definitions and axioms; otherwise they are quite inaccurate and insufferable, for they are only right when the principles are quite clear,” Blaise Pascal (1623 - 1664, French mathematician, author, and Catholic philosopher)